Atoms can bond to other atoms to make molecules. A
Answer #1

a) 1 m3 = 264 gallons

1 gallon = (1/264) m3

Density = 720 kg/m3

So, this means that mass of 1 m3 = 720 kg

So, mass of (1/264) m3 = (720/264) = 2.73 kg

Mass of 1 molecule = 18*10-26 kg

Number of molecules in 2.73 kg = 2.73/(18*10-26) = 0.15*1026 molecules = 15*1024 molecules

1 gallon has 15*1024 molecules.

b) 1 molecule of gasoline produces = 5*10-19 J

1 gallon has 15*1024 molecules.

So, 15*1024 molecules will produce energy = 5*10-19 J * 15*1024 = 75*105 J

1 gallon will produce 75*105 J.

c) mass = m = 1000 kg

speed = 29 m/s

Kinetic energy = (1/2)mv2 = (1/2)*1000*(29)2 = 420500 J = 4.20*105 J

d) 1 gallon will produce 75*105 J.

Kinetic energy = 4.20*105 J

So, gallons required to produce 4.20*105 J = (4.20*105)/(75*105) = 0.056 gallons

Gallons required = 0.056 gallon

e) 1 gallon = $ 2.20

Cost of 0.056 gallon = 0.056*2.20 = $ 0.12

The gasoline will cost $ 0.12.